# Extremal Process of Last Progeny Modified Branching Random Walks

### v1

@@ 1 Introduction
i.e. $\theta=\theta_{0}$ and that (1.5) holds. Let $\nu$ be a probability distribution on $\mathbb{R}$ such that there exists a regularly varying function $L$ at $\infty$ with index $\alpha\in(-2,0)$ satisfying (1.8). Then, writing
$m_{n}=n\frac{\kappa(\theta)}{\theta}+\frac{1}{\theta}\log L\left(\sqrt{n}\right)-\frac{1}{2\theta}\log n\\ \text{and}\quad c_{2}=\sqrt{\frac{2}{\pi\kappa^{\prime\prime}(\theta)}}\left(2\kappa^{\prime\prime}(\theta)\right)^{\frac{\alpha}{2}}\Gamma\left(\frac{\alpha}{2}+1\right),$
the extremal process $\tau_{-m_{n}}\mathcal{E}_{n}$ converges in law to a PPP$(c_{2}\theta Z_{\infty}e^{-\theta x}\,\mathrm{d}x)$.
@@
Theorem 1.4 also holds assuming that $\nu$ satisfies (1.8) with $L$ a positive constant, using the same arguments as the one we use below. This result therefore extends the previous estimate of Bandyopadhyay and Ghosh [6] for $\nu$ a Gumbel distribution with parameter $\frac{1}{\theta_{0}}$.
@@
We believe that up to adding a stronger integrability condition on the reproduction law of the BRW, one could prove a result similar to Theorem 1.4 assuming that $\nu$ satisfies (1.8) with $L$ a regularly varying function with parameter $\alpha\geq 0$. However, we predict a sharp phase correction around $\alpha=-2$, and that the limiting extremal process should resemble the one obtained in forthcoming Theorem 1.7 if $\alpha<-2$.
Finally, if $\nu$ satisfies (1.8) with $\theta>\theta_{0}$, then the asymptotic behaviour of $\mathcal{E}_{n}$ can be deduced from the asymptotic behaviour of $\mathcal{Z}_{n}$, the extremal process of the BRW.
@@
$\nu([x,\infty))\leq Ce^{-\theta x}\text{ for $x\in\mathbb{R}$.}$ (1.9)
@@
Theorem 1.7 follows as a direct consequence of the convergence in distribution of the extremal process of the branching random walk of Madaule [16]. The condition (1.9) is used to ensure that the point process defined in (1.10) is well-defined.
The rest of the article is organized as follows. We introduce in the next section some estimates that allow us to study the Laplace transform of $\mathcal{E}_{n}$. We then use the convergence of this Laplace transform to prove our main theorems in Section 3.
2 Laplace transform of the last progeny modified branching random walk
To prove the convergence in distribution of $\tau_{-m_{n}}\mathcal{E}_{n}$ to a limiting point measure $\mathcal{Z}$ in law for the topology of vague convergence, it is sufficient to show that for all continuous compactly supported function $\varphi$, we have
$\lim_{n\to\infty}{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}={\langle\mathcal{Z},\varphi\rangle}\quad\text{ in law,}$,
@@ 2 Laplace transform of the last progeny modified branching random walk
$\varphi\in\mathcal{T}\mapsto\operatorname{\mathbf{E}}\left(\exp\left(-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}\right)\right),$
@@
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left(\exp\left(-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}\right)\right)=\operatorname{\mathbf{E}}\left(\exp\left(-{\langle\mathcal{Z},\varphi\rangle}\right)\right).$ (2.1)
@@
$\begin{split}\operatorname{\mathbf{E}}\left(\exp\left(-{\langle\tau_{x}\mathcal{E}_{n},\varphi\rangle}\right)\right)&=\operatorname{\mathbf{E}}\left(\exp\left(-\sum_{|u|=n}g_{\varphi}(x+S_{u})\right)\right)\\ &=\operatorname{\mathbf{E}}\left(\exp\left(-{\langle\tau_{x}\mathcal{X}_{n},g_{\varphi}\rangle}\right)\right)\end{split}$ (2.2)
@@
$g_{\varphi}:x\mapsto-\log\int e^{-\varphi(x+y)}\,\nu(\mathrm{d}y),$
@@
As mentioned above, applying Madaule’s theorem [16] to (2.2) immediately yields Theorem 1.7. To prove Theorems 1.11.3 and 1.4, we first show that $\max_{|u|=n}S_{u}-m_{n}$ converges to $\infty$- ∞ in probability (for $m_{n}$ the process defined in each of these theorems), then study the asymptotic behaviour of $g_{\varphi}(x)$ as $x\to-\infty$. Let us first state the following generic lemma for BRW.
@@
If $\kappa(\theta)<\infty$, then $\lim_{n\to\infty}\max_{|u|=n}\theta S_{u}-n\kappa(\theta)=-\infty$ a.s.
Remark 2.2.
Recall that $W_{n}(\theta)=\sum_{|u|=n}e^{\theta S_{u}-n\kappa(\theta)}$ is a non-negative martingale, hence converges almost surely. As a result, using the bound
$W_{n}(\theta)\geq e^{\theta M_{n}-n\kappa(\theta)},$
it is immediate that $\sup_{n\in\mathbb{N}}\theta M_{n}-n\kappa(\theta)<\infty$ a.s.
Proof.
To prove that $\theta M_{n}-n\kappa(\theta)\to-\infty$, we strengthen the bound of Remark 2.2 with the following observation. The martingale $(W(\theta))$ converges almost surely, therefore $(W_{n}(\theta))$ is almost surely a Cauchy sequence.
Let $x\in\mathbb{R}$, we set $A_{x}:=\{\limsup_{n\to\infty}\theta M_{n}-n\kappa(\theta)>x\}$. We define a sequence $(T_{n})$ of stopping times by $T_{0}=0$ and
$T_{n+1}=\inf\{k>T_{n}:\theta M_{k}-k\kappa(\theta)>x\}.$
By definition, $T_{n}$ is finite for all $n$ on $A_{x}$. We define $A_{x,n}:=\{T_{n}<\infty\}$. Note that $A_{x,n}\in\mathcal{F}_{T_{n}}$ and $A_{x,n+1}\subseteq A_{x,n}$ for all $n$, and $A_{x}=\bigcap_{n\in\mathbb{N}}A_{x,n}$, where $(\mathcal{F}_{n})$ is the usual filtration of the BRW. We write $r_{n}$ for the label of one individual at generation $T_{n}$ such that $S_{r_{n}}=M_{T_{n}}$, and we define
$\overline{W}_{T_{n}+1}(\theta)=\sum_{|u|=T_{n}+1}e^{\theta S_{u}-(T_{n}+1)\kappa(\theta)}\mathbbm{1}_{\left\{u\not\succ r_{n}\right\}}+e^{\theta S_{r_{n}}-T_{n}\kappa(\theta)}\sum_{j\in\mathbb{N}}e^{\theta X_{j}-\kappa(\theta)},$
where $u\succ v$ means that $u$ is a descendant of $v$, and $(X_{j},j\in\mathbb{N})$ is a random sequence with same law as $(S_{u},|u|=1)$, independent of the BRW. In words, we obtain $\overline{W}_{T_{n}+1}(\theta)$ by replacing the offspring of an individual alive at position $M_{T_{n}}$ by an independent copy of that point process. It is therefore apparent that $\overline{W}_{T_{n}+1}(\theta)$ has, conditionally on $\mathcal{F}_{T_{n}}$, the same law as $W_{T_{n}+1}(\theta)$. Moreover, we have that on $A_{x,n}$,
$\left|W_{T_{n}+1}(\theta)-\overline{W}_{T_{n}+1}(\theta)\right|$ $=e^{\theta S_{r_{n}}-(T_{n}+1)\kappa(\theta)}\left|\sum_{j\in\mathbb{N}}e^{\theta(S_{r_{n}j}-S_{r_{n}})}-e^{\theta X_{j}}\right|$
$\geq e^{x-\kappa(\theta)}\left|\sum_{j\in\mathbb{N}}e^{\theta(S_{r_{n}j}-S_{r_{n}})}-e^{\theta X_{j}}\right|,$
with $r_{n}j$ being the $j$th descendant of individual $r_{n}$. In particular, there exists $\delta>0$ such that
$\operatorname{\mathbf{E}}\left(\mathbbm{1}_{A_{x,n}}\left(\left|W_{T_{n}+1}(\theta)-\overline{W}_{T_{n}+1}(\theta)\right|\wedge 1\right)\right)\geq\delta\operatorname{\mathbf{P}}(A_{x,n})\geq\delta\operatorname{\mathbf{P}}(A_{x}).$
On the other hand, we have
$\operatorname{\mathbf{E}}\left(\mathbbm{1}_{A_{x,n}}\left(\left|W_{T_{n}+1}(\theta)-W_{T_{n}}(\theta)\right|\wedge 1\right)\right)=\operatorname{\mathbf{E}}\left(\mathbbm{1}_{A_{x,n}}\left|\overline{W}_{T_{n}+1}(\theta)-W_{T_{n}}(\theta)\right|\wedge 1\right).$
Therefore, we have
$\operatorname{\mathbf{E}}\left(\mathbbm{1}_{A_{x,n}}\left(\left|W_{T_{n}+1}(\theta)-W_{T_{n}}(\theta)\right|\wedge 1\right)\right)\geq\delta\operatorname{\mathbf{P}}(A_{x})/2$
for all $n\in\mathbb{N}$. As $(W_{n}(\theta))$ converges almost surely as $n\to\infty$, we conclude by dominated convergence theorem that $\operatorname{\mathbf{P}}(A_{x})=0$.
To complete the proof, we observe that
$\operatorname{\mathbf{P}}(\limsup_{n\to\infty}\theta M_{n}-n\kappa(\theta)>-\infty)=\operatorname{\mathbf{P}}(\cup_{k\in\mathbb{N}}\operatorname{\mathbf{P}}(A_{-k}))=0.\qed$
Lemma 2.1 is enough to conclude that $\max_{|u|=n}S_{u}-m_{n}\to-\infty$ under the assumptions of Theorems 1.1 and 1.3. To obtain a similar result under the assumptions of Theorem 1.4, we will use the tightness of
@@
proved under the assumptions [^1] By [2, Equation (B1)], if $\kappa^{\prime\prime}(\theta_{0})<\infty$ then (1.5) implies [18, Equation (1.4)]. of Theorem 1.4 in [18]. Using the above observations, the asymptotic behaviour of the Laplace transform of $\mathcal{E}_{n}$ as $n\to\infty$ can be obtained by studying the asymptotic behaviour of $g_{\varphi}$ under assumption (1.8) as $x\to-\infty$.
@@
$\lim_{x\to-\infty}\frac{e^{-\theta x}}{L(-x)}g_{\varphi}(x)=\int\theta e^{-\theta z}(1-e^{-\varphi(z)})\,\mathrm{d}z.$
@@
$g_{\varphi}(x)=-\log\left(1-\int 1-e^{-\varphi(z+x)}\,\nu(\mathrm{d}z)\right),$
and that $y\mapsto 1-e^{-\varphi(y)}$ is continuous and compactly supported, i.e. an element of $\mathcal{T}$. By the dominated convergence theorem, we have that
$\lim_{x\to-\infty}\int 1-e^{-\varphi(z+x)}\,\nu(\mathrm{d}z)=0,$
@@
$g_{\varphi}(x)\sim\int 1-e^{-\varphi(z+x)}\,\nu(\mathrm{d}z)\quad\text{ as $x\to-\infty$.}$
@@
$\lim_{x\to-\infty}\int\frac{e^{-\theta x}}{L(-x)}\psi(x+z)\,\nu(\mathrm{d}z)=\int\theta e^{-\theta z}\psi(z)\,\mathrm{d}z$ (2.3)
for all $\psi\in\mathcal{T}$ to prove Lemma 2.3. We observe that for all $a<b$,
$\lim_{x\to-\infty}\frac{e^{-\theta x}}{L(-x)}\nu([a-x,b-x))=e^{-\theta a}-e^{-\theta b}=\int_{a}^{b}\theta e^{-\theta z}\,\mathrm{d}z,$
from which we deduce (2.3) by approximations, which completes the proof. ∎
@@
$c_{\varphi}(\theta)=\int\theta e^{-\theta z}(1-e^{-\varphi(z)})\,\mathrm{d}z.$
We can then restate Lemma 2.3 as $g_{\varphi}(x)\sim c_{\varphi}(\theta)L(-x)e^{\theta x}$ as $x\to-\infty$ under assumption (1.8).
@@
$g_{\varphi}(x)\leq C^{\prime}e^{\theta x}.$
@@
Let $\varphi\in\mathcal{T}$, as $\varphi$ is compactly supported, there exists $B>0$ such that $\varphi(z)=0$ for all $z<-B$. Therefore,
$e^{-g_{\varphi}(x)}=\int e^{-\varphi(x+z)}\,\nu(\mathrm{d}z)\geq\nu((-\infty,-x-B))\geq 1-Ce^{\theta(x+B)}.$
@@
$\limsup_{x\to-\infty}e^{-\theta x}g_{\varphi}(x)\leq\lim_{x\to-\infty}-e^{-\theta x}\log\left(1-Ce^{\theta(x+B)}\right)=Ce^{\theta B}.$
@@
and to identify its limit. This is mainly done using the so-called many-to-one lemma (see [19, Theorem 1.1]), that we now state.
@@
$\operatorname{\mathbf{E}}\left(\sum_{|u|=n}e^{\theta S_{u}-n\kappa(\theta)}f(S_{u_{1}},\ldots,S_{u_{n}})\right)=\operatorname{\mathbf{E}}\left(f(T_{1},\ldots,T_{n})\right).$
Moreover, $\operatorname{\mathbf{E}}(T_{1})=\kappa^{\prime}(\theta)$ whenever this quantity is well-defined.
3 Proof of the theorems
We prove in this section our main theorems. We first consider the asymptotic behaviour of $\mathcal{E}_{n}$ below the boundary case, i.e. when $\theta<\theta_{0}$, assuming that this quantity is well-defined. We then turn to the proof of Theorem 1.4, i.e. assuming that $\theta=\theta_{0}$. Finally, we prove Theorem 1.7 in Section 3.3.
3.1 BRW below the boundary case

### v2

@@ 1.2. Extremal process in the critical regime
i.e., $\theta=\theta_{0}$ and that (1.10) holds. Let $\nu$ be a probability distribution on $\mathbb{R}$ such that there exists a regularly varying function $L$ at $\infty$ with index $\alpha\in(-2,0)$ satisfying (1.5). Then, writing
$m_{n}=n\frac{\kappa(\theta)}{\theta}+\frac{1}{\theta}\log L\left(\sqrt{n}\right)-\frac{1}{2\theta}\log n\quad\text{and}\quad c_{2}=\sqrt{\frac{2}{\pi\kappa^{\prime\prime}(\theta)}}\left(2\kappa^{\prime\prime}(\theta)\right)^{\frac{\alpha}{2}}\Gamma\left(\frac{\alpha}{2}+1\right),$
[…]
@@
Theorem 1.4 also holds assuming that $\nu$ satisfies (1.5) with $L$ a positive constant, using the same proof techniques as the ones we develop below. This result therefore extends the previous estimate of Bandyopadhyay and Ghosh (2025) for $\nu$ a Gumbel distribution with parameter $\frac{1}{\theta_{0}}$.
@@
We believe that up to adding a stronger integrability condition on the reproduction law of the BRW, one could prove a result similar to Theorem 1.4 assuming that $\nu$ satisfies (1.5) with $L$ a regularly varying function with parameter $\alpha\geq 0$. However, we predict a sharp phase correction around $\alpha=-2$, and that the limiting extremal process should resemble the one obtained in the supercritical regime if $\alpha<-2$.
1.3. Extremal process in the supercritical regime
We finally turn to the case of perturbation law $\nu$ satisfying (1.5) with $\theta>\theta_{0}$. In this situation, the perturbation has such strong tails that its asymptotic behaviour can be deduced from the asymptotic behaviour of $\mathcal{Z}_{n}$, the extremal process of the BRW. As a result, we consider here a BRW satisfying the assumptions used by Madaule (2017) to prove the convergence of $\mathcal{Z}_{n}$. These conditions include notably (1.10). In this situation, it will be sufficient for the tail of $\nu$ to be light enough for the convergence of $\mathcal{E}_{n}$ to hold.
@@ 1.3. Extremal process in the supercritical regime
$\nu([x,\infty))\leq C\mathrm{e}^{-\theta x}\text{ for $x\in\mathbb{R}$.}$ (1.11)
@@
1.4. Heuristics for our results
The asymptotic behaviour of $\mathcal{E}_{n}$ can be explained by observing the distribution of particles at the $n$th generation of the BRW. We assume here that $\kappa(\theta)$ is finite for all $\theta\geq 0$, and that there exists $\theta_{0}$ satisfying (1.3). In this situation, for all $\theta<\theta_{0}$ it is well-known that the contributions to the value of the martingale $W_{n}(\theta)$ are mostly supported by particles in a neighbourhood of $n\kappa^{\prime}(\theta)$. Otherwise stated, there are at time $n$ approximately $W_{\infty}(\theta)\mathrm{e}^{-n(\theta\kappa^{\prime}(\theta)-\kappa(\theta))}$ particles to the neighbourhood of $n\kappa^{\prime}(\theta)$. Moreover, with high probability, there is no particle to the right of $n\kappa^{\prime}(\theta_{0})$.
Let us now estimate the maximal position reached by $S_{u}+Y_{u}$ by a particle $u$ such that $S_{u}$ is close to $n\kappa^{\prime}(\theta)$. Under the assumption that $\nu$ has an exponential tail with parameter $\vartheta$, we know that the maximum of $N$ independent copies of $Y$ will be close to $\frac{\log N}{\vartheta}$. Therefore, we expect particles in a neighbourhood of $n\kappa^{\prime}(\theta)$ to create a cloud of perturbed observations with the furthest point being found around
$n\left[\kappa^{\prime}(\theta)+\frac{1}{\vartheta}\left(\kappa(\theta)-\theta\kappa^{\prime}(\theta)\right)\right]+\frac{1}{\vartheta}\log W_{\infty}(\theta).$
Differentiating the linearly growing term with respect to $\theta$, we obtain $\kappa^{\prime\prime}(\theta)(1-\theta/\vartheta)$, therefore this quantity will be maximal, for large $n$, when $\theta=\min(\vartheta,\theta_{0})$.
When $\vartheta<\theta_{0}$, the extremal process $\mathcal{E}_{n}$ is thus obtained by generating a large number of i.i.d. copies of $Y$. As a result, this extremal process, centred by $n\kappa(\vartheta)/\vartheta$ will converge to a Poisson point process with intensity $\vartheta\mathrm{e}^{-\vartheta x}\,\mathrm{d}x$, shifted by $\frac{1}{\vartheta}\log W_{\infty}(\vartheta)$ to take into account the exact number of random variables. This corresponds to the situation described in Theorems 1.2 and 1.3.
When $\vartheta\geq\theta_{0}$, then the maximum of the last progeny modified BRW is generated by one of the leading particles, in a neighbourhood of $n\kappa^{\prime}(\theta_{0})$. This explains the shape of the limiting extremal process in the supercritical regime, which is obtained by perturbing the limiting extremal process of the BRW. In Theorem 1.7, the tail of the perturbations is so thin that only particles very close to the edge create a large perturbation, which yields the natural form of the limiting extremal process, obtained by perturbing the limiting extremal process of the original BRW with i.i.d. random variables. In Theorem 1.4, the main contribution near the tip of $\mathcal{E}_{n}$ comes from particles at distance $\Theta(\sqrt{n})$ from the leading edge of the BRW, which are counted using the derivative martingale of the BRW. This explains the extremal process, given by a Poisson point process with exponential intensity, shifted by $\frac{1}{\theta_{0}}\log Z_{\infty}$ instead of $\frac{1}{\theta_{0}}\log W_{\infty}(\theta_{0})$, and the apparition of $L(\sqrt{n})$ in the centring term $m_{n}$, as the typical perturbations observed at generation $n$ make jumps of that size $z\in\Theta(\sqrt{n})$, with probability approximately $L(z)\mathrm{e}^{-z}$.
Our proof techniques consist in validating this heuristic computation, by identifying the set of particles in the BRW that contribute to the generation of the tip of the point measure $\mathcal{E}_{n}$, then carefully proving the convergence towards the identified extremal process.
1.5. Organization of the article.
We introduce in the next section some estimates that allow us to study the Laplace transform of $\mathcal{E}_{n}$. We then use the convergence of this Laplace transform to prove our main theorems in Section 3.
2. Laplace transform of the last progeny modified branching random walk
To prove the convergence in distribution of $\tau_{-m_{n}}\mathcal{E}_{n}$ to a limiting point measure $\mathcal{Z}$ in law for the topology of vague convergence, it is sufficient to show that for any continuous compactly supported function $\varphi$, we have
$\lim_{n\to\infty}{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}={\langle\mathcal{Z},\varphi\rangle}\quad\text{ in law,}$
@@ 2. Laplace transform of the last progeny modified branching random walk
$\varphi\in\mathcal{T}\mapsto\operatorname{\mathbf{E}}\left[\exp\left(-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}\right)\right],$
@@
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left[\exp\left(-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}\right)\right]=\operatorname{\mathbf{E}}\left[\exp\left(-{\langle\mathcal{Z},\varphi\rangle}\right)\right].$ (2.1)
@@
$\operatorname{\mathbf{E}}\left[\exp\left(-{\langle\tau_{x}\mathcal{E}_{n},\varphi\rangle}\right)\right]=\operatorname{\mathbf{E}}\left[\exp\left(-\sum_{|u|=n}g_{\varphi}(x+S_{u})\right)\right]=\operatorname{\mathbf{E}}\left[\exp\left(-{\langle\tau_{x}\mathcal{X}_{n},g_{\varphi}\rangle}\right)\right]$ (2.2)
@@
$g_{\varphi}:x\mapsto-\log\int\mathrm{e}^{-\varphi(x+y)}\,\nu(\mathrm{d}y),$
@@
Applying Madaule’s theorem (Madaule, 2017) to (2.2) yields Theorem 1.7 (see Section 3.3). To prove Theorems 1.21.3, and 1.4, we first show that $M_{n}-m_{n}$ converges to -$\infty$ in probability, where we recall that $M_{n}:=\max_{|u|=n}S_{u}$ (with $m_{n}$ as defined in each theorem), and then study the asymptotic behaviour of $g_{\varphi}(x)$ as $x\to$-$\infty$. Let us first state the following generic lemma for BRW, which follows from (Shi, 2015, Lemma 3.1), using the transformation $x\mapsto-\theta x+\kappa(\theta)$ on the underlying point process.
@@
If $\kappa(\theta)<\infty$, then $\lim_{n\to\infty}\theta M_{n}-n\kappa(\theta)=$-$\infty$ a.s.
Lemma 2.1 is enough to conclude that $M_{n}-m_{n}\to$-$\infty$ under the assumptions of Theorems 1.2 and 1.3. To obtain a similar result under the assumptions of Theorem 1.4, we use the tightness of
@@
which was proved in Mallein (2016) under the assumptions [^2] By (Aïdékon, 2013, Equation (B.1)), if $\kappa^{\prime\prime}(\theta_{0})<\infty$ then (1.10) implies (Mallein, 2016, Equation (1.4)). of Theorem 1.4. Using the above observations, the asymptotic behaviour of the Laplace transform of $\mathcal{E}_{n}$ as $n\to\infty$ can be obtained by studying the asymptotic behaviour of $g_{\varphi}$ under assumption (1.5) as $x\to$-$\infty$.
@@
$\lim_{x\to-\infty}\frac{\mathrm{e}^{-\theta x}}{L(-x)}g_{\varphi}(x)=\int\theta\mathrm{e}^{-\theta z}(1-\mathrm{e}^{-\varphi(z)})\,\mathrm{d}z.$
@@
$g_{\varphi}(x)=-\log\left(1-\int 1-\mathrm{e}^{-\varphi(z+x)}\,\nu(\mathrm{d}z)\right),$
and that $y\mapsto 1-\mathrm{e}^{-\varphi(y)}$ is continuous and compactly supported, i.e., an element of $\mathcal{T}$. By the dominated convergence theorem, we have that
$\lim_{x\to-\infty}\int 1-\mathrm{e}^{-\varphi(z+x)}\,\nu(\mathrm{d}z)=0,$
@@
$g_{\varphi}(x)\sim\int 1-\mathrm{e}^{-\varphi(z+x)}\,\nu(\mathrm{d}z)\quad\text{ as $x\to-\infty$.}$
@@
$\lim_{x\to-\infty}\int\frac{\mathrm{e}^{-\theta x}}{L(-x)}\psi(x+z)\,\nu(\mathrm{d}z)=\int\theta\mathrm{e}^{-\theta z}\psi(z)\,\mathrm{d}z$ (2.3)
for all $\psi\in\mathcal{T}$ to prove Lemma 2.2. Using (1.5) together with the Uniform Convergence Theorem (Bingham et al., 1987, Theorem 1.5.2), we observe that for all $a<b$,
$\lim_{x\to-\infty}\frac{\mathrm{e}^{-\theta x}}{L(-x)}\nu([a-x,b-x))$
$\qquad=\lim_{x\to-\infty}\left(\mathrm{e}^{-\theta a}\frac{L(a-x)}{L(-x)}\frac{\mathrm{e}^{\theta(a-x)}}{L(a-x)}\nu([a-x,\infty))-\mathrm{e}^{-\theta b}\frac{L(b-x)}{L(-x)}\frac{\mathrm{e}^{\theta(b-x)}}{L(b-x)}\nu([b-x,\infty))\right)$
$\qquad=\mathrm{e}^{-\theta a}-\mathrm{e}^{-\theta b}=\int_{a}^{b}\theta\mathrm{e}^{-\theta z}\,\mathrm{d}z,$
from which we deduce (2.3) by approaching $\psi$ by a family of stair functions, thereby completing the proof. ∎
@@
$c_{\varphi}(\theta)=\int\theta\mathrm{e}^{-\theta z}(1-\mathrm{e}^{-\varphi(z)})\,\mathrm{d}z.$
We can then restate Lemma 2.2 as $g_{\varphi}(x)\sim c_{\varphi}(\theta)L(-x)\mathrm{e}^{\theta x}$ as $x\to$-$\infty$ under assumption (1.5).
@@
$g_{\varphi}(x)\leq C^{\prime}\mathrm{e}^{\theta x}.$
@@
Let $\varphi\in\mathcal{T}$, as $\varphi$ is compactly supported, there exists $B>0$ such that $\varphi(z)=0$ for all z<Bz<-B. Therefore,
$\mathrm{e}^{-g_{\varphi}(x)}=\int\mathrm{e}^{-\varphi(x+z)}\,\nu(\mathrm{d}z)\geq\nu((-\infty,-x-B))\geq 1-C\mathrm{e}^{\theta(x+B)}.$
@@
$\limsup_{x\to-\infty}\mathrm{e}^{-\theta x}g_{\varphi}(x)\leq\lim_{x\to-\infty}-\mathrm{e}^{-\theta x}\log\left(1-C\mathrm{e}^{\theta(x+B)}\right)=C\mathrm{e}^{\theta B}.$
@@
and to identify its limit. This is mainly done using the so-called many-to-one lemma (see Shi, 2015, Theorem 1.1), that we now state.
@@
$\operatorname{\mathbf{E}}\left[\sum_{|u|=n}\mathrm{e}^{\theta S_{u}-n\kappa(\theta)}f(S_{u_{1}},\ldots,S_{u_{n}})\right]=\operatorname{\mathbf{E}}\left[f(T_{1},\ldots,T_{n})\right].$
Moreover, $\operatorname{\mathbf{E}}[T_{1}]=\kappa^{\prime}(\theta)$ whenever this quantity is well-defined.
3. Proof of the theorems
We prove in this section our main theorems. We first consider the asymptotic behaviour of $\mathcal{E}_{n}$ below the boundary case, i.e., when $\theta<\theta_{0}$, assuming that this quantity is well-defined. We then turn to the proof of Theorem 1.4, i.e., assuming that $\theta=\theta_{0}$. Finally, we prove Theorem 1.7 in Section 3.3.
3.1. BRW below the boundary case